A rectangular cutting board is 6 inches longer than it is wide. If the area of the cutting board is 216 square inches, find the length and the width.
Answer by checkley77(12844)

L=W+6
LW=216
W(W+6)=216
W^2+6W=216
W^2+6W-216=0
(W-12)(W+18)=0
W-12=0
W=12 ANS. FOR THE WIDTH.
L=12+6=18 ANS. FOR THE LENGTH.
12*18=216
216=216